# Equivariant Flow Matching for Symmetry-Breaking Bifurcation Problems

### v1

@@ 4 Node Graph:
For this problem, we consider a beam that is attached to a base on the bottom, and then prescribe a displacement to the top of the beam. For a small displacement, the beam will compress but stay straight, but when the displacement reaches a critical value, the system becomes unstable and the beam will buckle, which will make it bend to either the left side or right side; either option is equally likely. We model the beam as a series of $n$ connected segments, which connect $n+1$ nodes, where node 0 is attached to the base. We consider $n$ ranging from 2 to 10. Each segment has a length $L_{i}$ and a stiffness $C_{i}$. At each node except for the top one the beam can bend, and therefore has a rotational stiffness $K_{i}$. The input $d$ is the vertical displacement applied to the top node. We solve the deformation of the beam by finding the Hencky strains $\epsilon_{i}$ and angles $q_{i}$ that minimize the strain energy density
@@
The Allen-Cahn equation is reaction-diffusion equation and describes the time evolution of a binary mixture of phases, e.g. ice and water at zero degrees Celsius. It is given by
@@ Appendix B Architectures & Training
For the VAE, we used an MLP with 2 hidden layers of 64 neurons per layer as an encoder and another as decoder, with a latent dimension of 16. For the flow model, we used an MLP with 3 hidden layers, 64 neurons per layer. Both Flow and VAE were trained for 10000 epochs with a batch size of 256. For the flow model we used a learning rate of 1e-3 and for the VAE a learning rate of 1e-2.
@@
As a non-probabilistic model we used an MLP with 2 hidden layers, with each 32 neurons per layer. It was trained for 52 epochs (determined by early stopping), with a batch size of 64 and a learning rate of 1e-4. For the VAE, we used an MLP with 2 hidden layers of 32 neurons as an encoder and a decoder, and a latent dimension of 16. We use another similar MLP to condition on the betted amount. The VAE was trained for 1000 epochs with a learning rate of 1e-3. For the flow model, we used a similar MLP to predict the flow field, and as a prior we used Gaussian noise with standard deviation 1.0.
@@ 4 Node Graph:
For the buckling beam, we used a random walk prior similar to the approach from Brinke et al.Brinke et al. [2025]. We used an EGNN Satorras et al. [2021] to predict interactions between nodes and a UNet for the time evolution.
@@
We used a random walk prior, which consists of cumulative, scaled random Gaussian noise. We then use a 2D UNet to predict the flow field, which maps an entire random walk trajectory to the entire final trajectory. It is periodic in only the spatial dimension. UNet has 4 downsampling steps and 32 channels.
@@ Appendix C Additional Results
Figure 6: Comparison of predicted Allen-Cahn trajectories by the flow model with examples of possible ground truth trajectories.

### v2

@@ 4 Node Graph:
For this problem, we consider a beam that is attached to a base on the bottom, and then prescribe a vertical displacement to the top of the beam, while leaving the horizontal displacement free. For a small displacement, the beam will compress but stay straight, but when the displacement reaches a critical value, the system becomes unstable and the beam will buckle, which will make it bend to either the left or right side, with either option equally likely. We model the beam as a series of $n$ connected segments, which connect $n+1$ nodes, where node 0 is attached to the base. We consider $n$ ranging from 2 to 10. Each segment has a length $L_{i}$ and a stiffness $C_{i}$. At each node except for the top one the beam can bend, and therefore has a rotational stiffness $K_{i}$. The input $d$ is the vertical displacement applied to the top node. We solve the deformation of the beam by finding the Hencky strains $\epsilon_{i}$ and angles $q_{i}$ that minimize the total energy:
@@
The Allen-Cahn equation is a reaction-diffusion equation and describes the time evolution of a binary mixture of phases, e.g., ice and water at zero degrees Celsius. It is given by
@@ Appendix B Architectures & Training
For the VAE, we used an MLP with 2 hidden layers of 64 neurons per layer as an encoder and another one as a decoder, with a latent dimension of 16. For the flow model, we used an MLP with 3 hidden layers, 64 neurons per layer. Both Flow and VAE were trained for 10000 epochs with a batch size of 256. For the flow model we used a learning rate of $10^{-3}$ and for the VAE a learning rate of $10^{-2}$.
@@
As a non-probabilistic model we used an MLP with 2 hidden layers, with each 32 neurons per layer. It was trained for 52 epochs (determined by early stopping), with a batch size of 64 and a learning rate of $10^{-4}$. For the VAE, we used an MLP with 2 hidden layers of 32 neurons as an encoder and a decoder, and a latent dimension of 16. We use another similar MLP to condition on the betted amount. The VAE was trained for 1000 epochs with a learning rate of $10^{-3}$. For the flow model, we used a similar MLP to predict the flow field, and as a prior we used Gaussian noise with standard deviation 1.0.
@@ 4 Node Graph:
For the buckling beam, we used a random walk prior similar to the approach from Brinke et al. brinke2025flow. We used an EGNN satorras2021n to predict interactions between nodes and a UNet for the time evolution.
@@
We used a random walk prior, which consists of cumulative, scaled random Gaussian noise. We then use a 2D UNet to predict the flow field, which maps an entire random walk trajectory to the entire final trajectory. It is periodic in only the spatial dimension. The UNet has 4 downsampling steps and 32 channels.
@@ Appendix C Additional Results
Figure 6: Comparison of predicted Allen-Cahn trajectories by the flow model with symmetric coupling with examples of possible ground truth trajectories.